# Observe Then Act: Asynchronous Active Vision-Action Model for Robotic Manipulation

### v1

@@ Abstract
In real-world scenarios, many robotic manipulation tasks are hindered by occlusions and limited fields of view, posing significant challenges for passive observation-based models that rely on fixed or wrist-mounted cameras. In this paper, we investigate the problem of robotic manipulation under limited visual observation and propose a task-driven asynchronous active vision-action model.Our model serially connects a camera Next-Best-View (NBV) policy with a gripper Next-Best Pose (NBP) policy, and trains them in a sensor-motor coordination framework using few-shot reinforcement learning. This approach allows the agent to adjust a third-person camera to actively observe the environment based on the task goal, and subsequently infer the appropriate manipulation actions. We trained and evaluated our model on 8 viewpoint-constrained tasks in RLBench. The results demonstrate that our model consistently outperforms baseline algorithms, showcasing its effectiveness in handling visual constraints in manipulation tasks. Our code and demonstration videos are available at https://hymwgk.github.io/ota/.
Index Terms:
Reinforcement learning , Robotic manipulation, Active vision, Next-best-view, Sensorimoter.
@@ I Introduction
Despite the reinforcement learning based models for robotic manipulation can learn to directly infer effective actions from camera-provided observation images through interaction with the environment, exhibiting strong scene perception and reasoning capabilities, they still heavily rely on cameras positioned based on prior knowledge, such as third-person fixed cameras and first-person wrist-mounted camera[1][2][3]. While these fixed third-person cameras and wrist-mounted cameras help reduce the sample variance of observations during interaction[3][4], they often struggle with providing sufficient visual information when dealing with manipulation scenarios under occlusions and limited observations due to unadjustable viewpoint positions or fixed camera-gripper configuration.
@@
Humans can independently adjust their head position and eye fixation based on their understanding of task goals and the current environment state to obtain the most advantageous visual information for making interactive action decisions[5]. Inspired by this, recent works in active vision-action joint learning[3][6][7][8] aim to mimic human capabilities by enabling agents to actively adjust their sensory “eye” viewpoints and interact with the world using their motor “hands.”
Similar to humans, the active vision-action model exists in two categories: synchronous and asynchronous[10]. The synchronous vision-action model makes real-time decisions on the next visual sensor and motor actions based on current observation information[7][8][11][12]. This setup requires the model to effectively coordinate sensor motor actions in the next moment and to infer motor actions even when sufficient observation information may not be available which impose higher demands on the model’s dynamic modeling capabilities, usually requiring more interaction steps and making it difficult to apply to manipulation tasks requiring large viewpoint changes.
Although humans are capable of making synchronous decisions, hand movements and changes in viewpoint typically do not occur simultaneously; instead, we generally select the optimal eye viewpoint before executing an interaction. Similarly, the asynchronous vision-action model separates visual sensing and motor actions over time, allowing the vision policy to focus less on coordinating both actions simultaneously when determining the observation viewpoint. Instead, the model can focus more on task-specific manipulations and distribute the sensor-motor coordination throughout the entire task cycle. Although such models are currently more often applied to object grasping tasks in occluded scenes[13][14], we believe that the asynchronous approach may be more suitable for common sequential manipulation tasks.
@@
In this paper, we further explore the problem of robotic manipulation under limited observation and introduce a task-driven asynchronous active vision-action model. The model is designed as a sequential dual-agent structure, as illustrated in Fig.LABEL:fig:head_image. It consists of two agents: a Next-Best-View (NBV) agent for inferring optimal camera viewpoints, and a Next-Best-Pose (NBP) agent for determining the gripper’s action based on the observations from the NBV-inferred viewpoint. By alternating between sensor and motor action inference in each episode, our model enables active perception of the environment and effective interaction with the environment to achieve the manipulation task objectives.
@@
We train our model using a few-shot reinforcement learning approach[2] and introduce a viewpoint-aware demo augmentation method to maximize the utilization of limited demonstration trajectories. To address the challenge posed by the expanded observation-action space resulting from the active viewpoint, we employ a viewpoint-centric voxel alignment technique, which rotates the scene voxel to align with the camera viewpoint. To further stabilize the training process, we incorporate a combined task-agnostic auxiliary reward which provides sufficiently task-aligned learning signals throughout the training process, encouraging the agent to interact effectively with the environment and select viewpoints that offer richer visual information for more informed gripper action decisions. We conduct comparison experiments and analyses on 8 manipulation tasks with limited effective viewpoints in RLbench[9]. The results demonstrate our model’s effectiveness in selecting task-relevant viewpoints and its robustness in inferring manipulation actions based on the current state and task requirements. The main contributions of this paper are as follows:
@@ II Related Work
Active Vision for Manipulation. Active vision aims to enable agents to actively observe the environment based on specific task goals. It has been widely applied in tasks such as indoor visual navigation[12][15], object searching[16, 17, 18, 19] and 3D reconstruction[20]. In robotic manipulation, active vision is primarily used for grasping pose detection in occluded environments[7][13][14][21][22] and for detecting fruit nodes in automated harvesting[23, 24, 25]. Recent studies have extended the use of active vision to more general robotic manipulation tasks. For instance, Cheng et al[11] proposed an active vision model with a target detector module to handle object-pushing task in occluded scenes. Lv et al[8] introduced a model-based sensing-aware framework for manipulating flexible objects. Matthew et al[4] employed a self-prediction approach to learn viewpoint adjustments while simultaneously developing motion skills for robotic arm joints. Similarly, Shang et al[6] utilized a self-supervised reward mechanism to jointly learn visual sensing and motor skills. Unlike these methods, our active vision-aciton model does not require a task-specific target detector and employs an RL-based dual agent asynchronous decision-making structure.
3D Voxel-based Manipulation. 3D voxels are highly effective in representing spatial structures[26], making them widely used in domains such as medical imaging[27][28] and autonomous driving[29][30]. Recently, they have also been adopted in robotic manipulation models for visual representation. For instance, James et al.[2] transformed original RGB-D images into 11-dimensional voxel grids and developed a 3DUnet-based reinforcement learning model to extract scene features and infer manipulation actions. Similarly, Shridhar et al.[31] used discrete 3D voxel images as input, employing the Perceiver network[32] as a backbone for end-to-end behavioral cloning, trained on demonstration trajectories. Leveraging the strong spatial structure representation and robust structural invariance of voxels under viewpoint changes—which aids in viewpoint learning—our model also uses voxelized scene images as input.
Figure 2: Comparison of regular RL and asynchronous active vision-action decision processes. We introduce an intermediate time step $t^{\prime}$ between $t$ and $t+1$, dividing the interaction into NBV policy for viewpoint selection and the NBP policy for action selection. The task reward $r_{\text{t'}}$ for state $\mathbf{s}$ is derived from the reward $r_{\text{t+1}}$ after the NBP interaction, which enables sensor-motor joint training through shared task rewards.
@@
Figure 3: Overview of our model pipeline. The NBV policy infers the scene ROI position $\mathbf{f}^{*}_{t}$ based on the current observation and determines the optmal viewpoint $\mathbf{v}^{*}$ for observing the ROI to achieve the task goal. The NBP agent infers the gripper actions based on the scene ROI image from viewpoint $\mathbf{v}^{*}$ and the proprioceptive state.
@@
Problem Formulation. We divide one asynchronous vision-action time step $t\sim t+1$ into the viewpoint decision and transform process of the NBV policy from $t\sim t^{\prime}$ and the gripper action decision and interaction process of the NBP policy from $t^{\prime}\sim t+1$. Therefore the asynchronous active vision manipulation problem can be formalized as a POMDP described by the tuple $<\mathcal{O},\mathcal{A^{\text{c}}},P,\mathcal{A^{\text{g}}},\mathcal{O^{\prime}},P^{\prime},R,\gamma>$ As shown in Fig.2, $\mathcal{O}$ and $\mathcal{O^{\prime}}$ represent the observation spaces of the agent at times $t$ and $t^{\prime}$, respectively. The NBV policy $\pi_{\text{v}}$ determines the camera viewpoint action $\mathbf{a}_{t}^{\text{c}}=\pi_{\text{v}}(\cdot|\mathbf{o}_{t})\in\mathcal{A^{\text{c}}}$ based on the observation $\mathbf{o}_{t}\in\mathcal{O}$, and obtains a new observation $\mathbf{o}_{t^{\prime}}\in\mathcal{O^{\prime}}$ through transition probability $p(\mathbf{o}_{t^{\prime}}|\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}})\in P$. The NBP policy $\pi_{\text{g}}$ then determines the gripper action $\mathbf{a}_{t^{\prime}}^{\text{g}}=\pi_{\text{p}}(\cdot|\mathbf{o}_{t^{\prime}})\in\mathcal{A^{\text{g}}}$ based on the new observation $\mathbf{o}_{t^{\prime}}$. Through the transition probability $p^{\prime}(\mathbf{o}_{t+1}|\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}})\in P^{\prime}$, scene observation $\mathbf{o}_{t+1}$ can be obtained by camera at time $t+1$, and the environment provides the task reward $r_{t+1}\in R$. Our goal is to jointly learn the policies $\pi^{*}_{\text{v}}$ and $\pi^{*}_{\text{g}}$ that maximize the expected sensor-motor joint reward for the given manipulation task, which is $\pi^{*}_{\text{v}},\pi^{*}_{\text{g}}=\arg\underset{\pi_{\text{v}},\pi_{\text{p}}}{\max}\mathbb{E}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}R(\mathbf{o}_{t},\mathbf{a}^{\text{c}}_{t},\mathbf{a}^{\text{g}}_{t^{\prime}})\Big{]}$.
@@
Observation and Action Space We define the observation spaces for the NBV and NBP policies as identical, with $\mathcal{O}=\mathcal{O^{\prime}}=\{\mathcal{O}^{\text{v}},\mathcal{O}^{\text{p}}\}$, where $\mathbf{o}^{\text{v}}\in\mathcal{O}^{\text{v}}$ represents the 3D voxel grid voxelized from the scene RGB-D image captured by the camera, and $\mathbf{o}^{\text{p}}\in\mathcal{O}^{\text{p}}$ is the proprioceptive state. Specifically, the raw RGB-D image $\mathbf{o}^{\text{i}}$ is converted into a point cloud $\mathbf{o}^{\text{pc}}\in\mathbb{R}^{n\times 6}$ using the function $\mathbf{P}(\mathbf{o}^{\text{i}}):\mathbf{o}^{\text{i}}\rightarrow\mathbf{o}^{\text{pc}}$, which is then voxelized through function $\mathbf{V}(\mathbf{o}^{\text{pc}},\mathbf{e},\mathbf{c}):\mathbf{o}^{\text{pc}}\rightarrow\mathbf{o}^{\text{v}}$, where $\mathbf{e}$ denotes the voxel resolution and $\mathbf{c}$ represents the voxel center coordinates. The resulting 3D voxel grid $\mathbf{o}^{\text{v}}\in\mathbb{R}^{n\times(3+M+1)}$ contains the center coordinates, $M$-dimensional features, and a occupancy flag for each voxel. The proprioceptive state $\mathbf{o}^{\text{p}}=\{\mathbf{v},\mathbf{g}\}$ comprises the camera viewpoint position $\mathbf{v}\in\mathbb{R}^{3}$ and the gripper pose $\mathbf{g}\in\mathbb{R}^{6}$.
@@
Following common settings[14][21][22], we restrict the camera viewpoint’s movement to the upper hemisphere at a fixed radius of $r^{\text{v}}=1.2\text{m}$ centered on the table, illustrated in Fig.8. The camera’s optical axis is assumed to always be directed towards the table center thus the viewpoint position can be described by spherical coordinates $\mathbf{v}=(r^{\text{v}},\theta^{\text{v}},\phi^{\text{v}})$. We than define the NBV policy’s action as $\mathbf{a}^{\text{c}}=\{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}\}\in\mathcal{A}^{\text{c}}$, where $\mathbf{k}^{\text{v}}\in\mathbb{R}^{2}$ represents the discrete viewpoint index for $(\theta^{\text{v}},\phi^{\text{v}})$, and $\mathbf{k}^{\text{f}}\in\mathbb{R}^{3}$ denotes the discrete center position index of a region of interest(ROI) that requires attention within the scene. The NBP policy’s action is defined as $\mathbf{a}^{\text{g}}=\{\mathbf{k}^{\text{g}},d\}\in\mathcal{A}^{\text{g}}$, where $\mathbf{k}^{\text{g}}$ represents the discrete 6-DoF gripper pose index, and $d$ is the goal distance between the parallel gripper tips.
@@ IV Method
In this section, we provide a detailed explanation of the fundamental structure of our active vision-action model, along with three innovative strategies designed to enhance both model performance and learning efficiency.
@@ IV-A Model Structure
Fig3 illustrates the basic pipline of our model. At time $t$, the voxelized image observation and proprioceptive state undergo a coordinate transformation for viewpoint alignment (Section IV-B), resulting in the processed observation $\mathbf{o}_{t}$. The NBV policy network, parameterized by $\Theta$, uses $\mathbf{o}_{t}$ to determine the discrete position index $\mathbf{k}^{\text{f}}$ of the ROI center in the scene, as described in Eq.(1). Simultaneously, the network infers the corresponding discrete index $\mathbf{k}^{\text{v}}$, representing the viewpoint that would best improve task success. The camera is then adjusted to the goal viewpoint $\mathbf{v}^{*}$, which is derived from the selected $\mathbf{k}^{\text{v}}$.
@@
Based on the predicted ROI position $\textbf{f}^{*}$, derived from $\mathbf{k}^{\text{f}}$, the raw image observation from the viewpoint is cropped and aligned to generate the corresponding ROI 3D voxel grid $\mathbf{o}^{\text{v}}_{t^{\prime}}$. This enables us to construct the observation $\mathbf{o}_{t^{\prime}}=\{\mathbf{o}^{\text{v}}_{t^{\prime}},\mathbf{o}^{\text{p}}_{t^{\prime}}\}$. Using the NBP policy network parameterized by $\Phi$, we infer the gripper closure action $d$ and the discrete gripper pose index $\mathbf{k}^{\text{g}}$ which is then transformed into the goal pose $\mathbf{g}^{*}$. Finally, a complete sensor-motor interaction step is achieved after gripper exectuation and rewarded by $r_{t+1}$.
@@
We denote the sub-transition for the NBV policy as $\text{T}_{t}^{\text{v}}=<\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}},\mathbf{o}_{t^{\prime}}>$ and for the NBP policy as $\text{T}_{t}^{\text{p}}=<\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}},\mathbf{o}_{t+1}>$. Since the NBV and NBP policies share the same reward $r_{t+1}$, a complete interaction transition is denoted as $\text{T}_{t}=\{\text{T}_{t}^{\text{v}},\text{T}_{t}^{\text{p}},r_{t+1}\}$. Upon completing an interaction episode, the collected transition is stored in the replay buffer together with the demonstration transitions. During model updates, both the NBV and NBP policies are updated using the same reward, thereby achieving sensor-motor joint learning. We employ a 3D-CNN based U-Net[35] as the image encoder for both policy networks, and three-layer fully connected network (FCN) for decoding the viewpoint and gripper actions.

### v2

@@ Abstract
In real-world scenarios, many robotic manipulation tasks are hindered by occlusions and limited fields of view, posing significant challenges for passive observation-based models that rely on fixed or wrist-mounted cameras. In this paper, we investigate the problem of robotic manipulation under limited visual observation and propose a task-driven asynchronous active vision-action model. Our model serially connects a camera Next-Best-View(NBV) policy with a gripper Next-Best-Pose(NBP) policy, and trains them in a sensor-motor coordination framework using few-shot reinforcement learning. This approach enables the agent to reposition a third-person camera to actively observe the environment based on the task goal, and subsequently determine the appropriate manipulation actions. We trained and evaluated our model on 8 viewpoint-constrained tasks in RLBench. The results demonstrate that our model consistently outperforms baseline algorithms, showcasing its effectiveness in handling visual constraints in manipulation tasks. Our code and demonstration videos are available at https://hymwgk.github.io/ota/.
@@ I INTRODUCTION
Although reinforcement learning-based models for robotic manipulation can learn to infer actions directly from image observations through interaction with the environment, they still heavily rely on carefully positioned static third-person or wrist-mounted cameras[1, 2, 3]. While these cameras reduce sample variance during interaction[3, 4], they often struggle with providing sufficient visual data in occluded or limited-visibility scenarios due to unadjustable viewpoints or fixed camera-gripper configuration.
@@
However, humans can adjust their head and eye positions based on task goals. This helps them obtain advantageous visual information for action decision [5]. Inspired by this, recent active vision-action models enable agents to actively adjust their sensory viewpoints while interacting with the environment[3, 6, 7, 8], mimicking human capabilities.
Active vision-action models, like human perception, can be categorized as synchronous or asynchronous[10]. Synchronous models make decisions about visual and motor actions simultaneously based on current observations[7, 8, 11, 12]. However, the need for simultaneous coordination of sensor-motor actions places high demands on the model’s dynamic modeling capabilities and may limits its ability to handle tasks requiring large viewpoint shifts.
In contrast, asynchronous models decouple vision sensing and motor actions over time, allowing the agent to prioritize observation before interaction and then focus on inferring actions based on the selected viewpoint, thereby distributing sensor-motor coordination across the entire task episode. Although such models are commonly used for object grasping in occluded scenes[13, 14], we believe that the asynchronous approach may also be suitable for sequential manipulation tasks.
@@
In this paper, we introduce a task-driven asynchronous active vision-action model for robotic manipulation under limited observation(Fig.LABEL:fig:head_image). Our model employs a sequential dual-agent structure, with a Next-Best-View(NBV) agent to infer optimal viewpoints and a Next-Best-Pose(NBP) agent to determine gripper actions based on the observations from the NBV-inferred viewpoints(Fig.3). By alternating between sensor and motor actions, the model actively perceives and interacts with the environment to achieve task objectives.
@@
We train our model with few-shot reinforcement learning[2] and introduce a viewpoint-aware demo augmentation method to maximize the utilization of limited demonstration trajectories. To address the expanded observation-action space caused by the active viewpoint, we employ a viewpoint-centric voxel alignment technique that rotates the scene grid to align with the camera viewpoint. We also integrate a task-agnostic auxiliary reward to stabilize training, encouraging agents to interact effectively and select informative viewpoints for gripper action decisions. We evaluate our model on 8 RLBench[9] tasks with limited viewpoints, demonstrating its effectiveness in viewpoint selection and action inference based on task requirements. The main contributions of this paper are as follows:
@@ II Related Work
Active Vision for Manipulation. Active vision aims to enable agents to actively observe the environment based on specific task goals. It has been widely applied in tasks such as indoor visual navigation[12, 15], object searching[16, 17] and 3D reconstruction[18]. In robotic manipulation, active vision is primarily used for grasping pose detection in occluded environments[7, 13, 14, 19] and for detecting fruit nodes in automated harvesting[20, 21, 22]. Recent studies have extended the use of active vision to more general robotic manipulation tasks. For instance, Cheng et al[11] proposed an active vision model with a target detector module to handle object-pushing task in occluded scenes. Lv et al[8] introduced a model-based sensing-aware framework for manipulating flexible objects. Matthew et al[4] employed a self-prediction approach to learn viewpoint adjustments while simultaneously developing motion skills for robotic arm joints. Similarly, Shang et al[6] utilized a self-supervised reward mechanism to jointly learn visual sensing and motor skills. Unlike these methods, our active vision-aciton model does not require a task-specific target detector and employs an RL-based dual agent asynchronous decision-making structure.
3D Voxel-based Manipulation. 3D voxels are highly effective in representing spatial structures[23], making them widely used in domains such as medical imaging[24, 25] and autonomous driving [26, 27]. Recently, they have also been adopted in robotic manipulation models for visual representation. For instance, James et al.[2] transformed original RGB-D images into 11-dimensional voxel grids and developed a 3DUnet-based reinforcement learning model to extract scene features and infer manipulation actions. Similarly, Shridhar et al.[28] used discrete 3D voxel images as input, employing the Perceiver network[29] as a backbone for end-to-end behavioral cloning, trained on demonstration trajectories. Leveraging the strong spatial structure representation and robust structural invariance of voxels under viewpoint changes—which aids in viewpoint learning—our model also uses voxelized scene images as input.
Figure 2: Decision process comparison.We divide single step interaction into NBV for viewpoint and NBP for action selection. Reward $r_{\text{t'}}$ is derived from $r_{\text{t+1}}$ after NBP interaction, enabling sensor-motor joint training through shared task rewards.
@@
Figure 3: Overview of our model pipeline. The NBV policy infers the scene ROI position $\mathbf{f}^{*}_{t}$ based on the current observation and determines the optimal viewpoint $\mathbf{v}^{*}$ for observing the ROI to achieve the task goal. The NBP agent infers the gripper actions based on the scene ROI image from viewpoint $\mathbf{v}^{*}$ and the proprioceptive state.
III Problem Formulation
@@
We divide one interaction time step $t\sim t+1$ into the viewpoint decision and transform process of the NBV policy $\pi_{\text{v}}$ from $t\sim t^{\prime}$ and the gripper action decision and interaction process of the NBP policy $\pi_{\text{p}}$ from $t^{\prime}\sim t+1$. The asynchronous active vision manipulation problem can be formalized as a POMDP described by the tuple $<\mathcal{O},\mathcal{A^{\text{c}}},P,\mathcal{A^{\text{g}}},P^{\prime},r,\gamma>$, as shown in Fig.2. where $\mathcal{O}$ is the observation space, $\mathcal{A^{\text{c}}},\mathcal{A^{\text{g}}}$ is the NBV and NBP action space, respectively. At time $t$, the NBV policy $\pi_{\text{v}}$ selects a camera action $\mathbf{a}_{t}^{\text{c}}=\pi_{\text{v}}(\cdot|\mathbf{o}_{t})\in\mathcal{A^{\text{c}}}$ based on the observation $\mathbf{o}_{t}\in\mathcal{O}$, resulting in a new observation $\mathbf{o}_{t^{\prime}}\in\mathcal{O^{\prime}}$ via transition probability $p(\mathbf{o}_{t^{\prime}}|\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}})\in P$. The NBP policy $\pi_{\text{p}}$ then selects a gripper action $\mathbf{a}_{t^{\prime}}^{\text{g}}=\pi_{\text{p}}(\cdot|\mathbf{o}_{t^{\prime}})\in\mathcal{A^{\text{g}}}$, with the next scene observation $\mathbf{o}_{t^{\prime}}$ obtained through $p^{\prime}(\mathbf{o}_{t+1}|\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}})\in P^{\prime}$. The environment provides a reward $r_{t+1}\in\mathbb{R}$. Our goal is to jointly learn $\pi^{*}_{\text{v}}$ and $\pi^{*}_{\text{p}}$ to maximize the expected sensor-motor reward $\mathbb{E}_{\pi_{\text{v}},\pi_{\text{p}}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r(\mathbf{o}_{t},\mathbf{a}^{\text{c}}_{t},\mathbf{a}^{\text{g}}_{t^{\prime}})\Big{]}$.
@@
Observation and Action Space. We define the observation spaces for the NBV and NBP policies as identical, with $\mathcal{O}=\{\mathcal{O}^{\text{v}},\mathcal{O}^{\text{p}}\}$, where $\mathbf{o}^{\text{v}}\in\mathcal{O}^{\text{v}}$ represents the 3D voxel grid derived from the RGB-D image, and $\mathbf{o}^{\text{p}}\in\mathcal{O}^{\text{p}}$ is the proprioceptive state. Specifically, the RGB-D image $\mathbf{o}^{\text{i}}$ is converted into a point cloud $\mathbf{o}^{\text{pc}}\in\mathbb{R}^{n\times 6}$ using the function $\mathbf{P}(\mathbf{o}^{\text{i}}):\mathbf{o}^{\text{i}}\rightarrow\mathbf{o}^{\text{pc}}$, and voxelized via $\mathbf{V}(\mathbf{o}^{\text{pc}},\mathbf{e},\mathbf{c}):\mathbf{o}^{\text{pc}}\rightarrow\mathbf{o}^{\text{v}}$, where $\mathbf{e}$ denotes the voxel resolution and $\mathbf{c}$ represents the voxel center. The resulting 3D voxel grid $\mathbf{o}^{\text{v}}\in\mathbb{R}^{n\times(3+M+1)}$ contains the center coordinates, $M$-dimensional features, and an occupancy flag for each voxel. The proprioceptive state $\mathbf{o}^{\text{p}}=\{\mathbf{v},\mathbf{g}\}$ comprises the camera viewpoint position $\mathbf{v}\in\mathbb{R}^{3}$ and the gripper pose $\mathbf{g}\in\mathbb{R}^{6}$.
@@
Following common practice[14, 19], the camera’s movement is limited to the upper hemisphere at a fixed radius $r^{\text{v}}=1.2\text{m}$ above the table, with the optical axis directed at the sphere center, illustrated in Fig.8(b). The viewpoint is described by spherical coordinates $\mathbf{v}=(r^{\text{v}},\theta^{\text{v}},\phi^{\text{v}})$. The NBV policy action is $\mathbf{a}^{\text{c}}=\{\mathbf{k}^{\text{v}},\mathbf{k}^{\text{f}}\}$, where $\mathbf{k}^{\text{v}}\in\mathbb{R}^{2}$ represents the discrete viewpoint index for $(\theta^{\text{v}},\phi^{\text{v}})$, and $\mathbf{k}^{\text{f}}\in\mathbb{R}^{3}$ denotes the discrete center position index of a region of interest(ROI) that requires attention within the scene. The NBP policy action $\mathbf{a}^{\text{g}}=\{\mathbf{k}^{\text{g}},d\}$ consists of the discrete 6-DoF gripper pose index $\mathbf{k}^{\text{g}}$ and the goal distance $d$ between the gripper tips.
@@ IV-A Model Structure
We illustrates the basic pipline of our model in Fig3. Once the alignment process is completed(Section IV-B), the observation $\mathbf{o}_{t}$ is obtained. The DQN-based NBV policy network, parameterized by $\Theta$, then uses $\mathbf{o}_{t}$ to determine the discrete position index $\mathbf{k}^{\text{f}}$ of the ROI center in the scene, as described in Eq.(1). Simultaneously, the network infers the corresponding discrete index $\mathbf{k}^{\text{v}}$, which represents the viewpoint most likely to enhance task success. The camera is subsequently adjusted to the goal viewpoint $\mathbf{v}^{*}$, derived from the selected $\mathbf{k}^{\text{v}}$.
@@
Based on the ROI position $\textbf{f}^{*}$, derived from $\mathbf{k}^{\text{f}}$, the image from $\mathbf{v}^{*}$ is cropped and aligned to generate the 3D voxel grid $\mathbf{o}^{\text{v}}_{t^{\prime}}$ for the ROI. The NBP policy network, also DQN-based and parameterized by $\Phi$, infers the gripper closure $d$ and pose index $\mathbf{k}^{\text{g}}$ based on $\mathbf{o}_{t^{\prime}}=\{\mathbf{o}^{\text{v}}_{t^{\prime}},\mathbf{o}^{\text{p}}_{t^{\prime}}\}$. Finally, a complete sensor-motor interaction step is achieved after gripper interaction and rewarded by $r_{t+1}$.
@@
We denote a complete interaction transition as $\text{T}_{t}=\{\text{T}_{t}^{\text{v}},\text{T}_{t}^{\text{p}},r_{t+1}\}$, where $\text{T}_{t}^{\text{v}}=(\mathbf{o}_{t},\mathbf{a}_{t}^{\text{c}},\mathbf{o}_{t^{\prime}})$ represents the NBV policy’s sub-transition, and $\text{T}_{t}^{\text{p}}=(\mathbf{o}_{t^{\prime}},\mathbf{a}_{t^{\prime}}^{\text{g}},\mathbf{o}_{t+1})$ represents the NBP policy’s sub-transition. After each training episode, transitions are stored in the replay buffer along with demonstration data. Both policies are updated using the same task reward $r_{t+1}$, thereby achieving sensor-motor joint learning. Each policy employs a separate 3D-CNN based U-Net[32] as the image encoder, followed by a three-layer fully connected network(FCN) for action decoding.